OpenStudy (anonymous):

Evaluate 9x^2y^-2 for x = –3 and y = 2. A. 324 B. 20/1/4 C. 9(-6)^0 D. 1/144

OpenStudy (anonymous):

81/2

OpenStudy (campbell_st):

\[9\times(-3)^2\times \frac{1}{2^2}\]

OpenStudy (campbell_st):

just evaluate for the correct answer

OpenStudy (anonymous):

the answer is B. 20/1/4 correct

OpenStudy (anonymous):

is this correct doubledrive?

OpenStudy (campbell_st):

thats correct.... doubledrive's answer is incorrect